Subject: [Fwd: Comment raith SNF 2009-04-23 09:09:04: Cooling loop flow check 9 AM 4-23-09]
From: James Conway <jwc@snf.stanford.edu>
Date: Thu, 23 Apr 2009 09:13:05 -0700
Thu, 23 Apr 2009 09:13:05 -0700
Good Morning Raith Users and Raith Field Service:

Good flow and pressure were observed this morning after repairs effected 
late yesterday to the cooling loop.
The Fe-Gun has returned to normal emission levels after its start up 
last evening.

Users are requested to call me immediately if any cooling or operations 
issues are encountered.

Thank you for your support!

James Conway


-------- Original Message --------
Subject: 	Comment raith SNF 2009-04-23 09:09:04: Cooling loop flow check 
9 AM 4-23-09
Date: 	Thu, 23 Apr 2009 09:09:05 -0700
From: 	jwc@snf.stanford.edu
To: 	raith-pcs@snf.stanford.edu



I am pleased to report that my efforts to pickle the cooling line loop to the raith
appears to have be successful.
After repairs were completed last evening the measured flow rate was 2.5 lpm at ~ 100
psi measured on the output side of the chiller.
This morning inspection measured 3.0 lpm plus and a very nice flow of clean solution
is circulating through the loop.
I am going to continue allowing this solution to circulate through the system over the
weekend until RAITH Field Service personnel are available to check the system out and
flush and refill with normal solutions.
We will monitor flow rate and pressure twice each working day, or if problems are encountered
on the tool.
JWC




Good Morning Raith Users and Raith Field Service:

Good flow and pressure were observed this morning after repairs effected late yesterday to the cooling loop.
The Fe-Gun has returned to normal emission levels after its start up last evening.

Users are requested to call me immediately if any cooling or operations issues are encountered.

Thank you for your support!

James Conway


-------- Original Message --------
Subject: Comment raith SNF 2009-04-23 09:09:04: Cooling loop flow check 9 AM 4-23-09
Date: Thu, 23 Apr 2009 09:09:05 -0700
From: jwc@snf.stanford.edu
To: raith-pcs@snf.stanford.edu


I am pleased to report that my efforts to pickle the cooling line loop to the raith appears to have be successful.
After repairs were completed last evening the measured flow rate was 2.5 lpm at ~ 100 psi measured on the output side of the chiller.
This morning inspection measured 3.0 lpm plus and a very nice flow of clean solution is circulating through the loop.
I am going to continue allowing this solution to circulate through the system over the weekend until RAITH Field Service personnel are available to check the system out and flush and refill with normal solutions.
We will monitor flow rate and pressure twice each working day, or if problems are encountered on the tool.
JWC